• In Young's double - slit experiment, a monochromatic parallel light of wavelength, λ=550 nm illuminates the double slits with separation of a=2×10⁻⁴m. The distance between the viewing screen and the barrier containing the double slits is D=2m. Find : (1) The distance between the pair of the tenth order bright fringes that are on both sides of the central maximum? (2) If one slit is covered by a glass sheet of thickness =6.6×10⁻⁵ m and index of refraction n=1.58, the zeroth maximum will move to a new position which corresponds to the original Kth order bright fringe. What is the value of K?
1
Expert's answer
2021-06-16T14:22:50-0400

Gives

D=2 m

"\mu" =1.58

"\lambda=500nm"

"d=2\times10^{-4} m"

(n)=10

"x=\frac{n\lambda D}{d}"

"x=\frac{10\times 550\times10^{-9} \times2}{2\times10^{-4}}"

x=0.055m

Part(b)

"t=6.6\times10^{-5}m"

"2(\mu -1)t=2x"


"x=(1.58-1)\times6.6\times10^{-5}=3.828\times10^{-5}m"

Find k Value

"x=\frac{k\lambda D}{d}"

Put value

"k=\frac{xd}{\lambda D}"

"k=\frac{3.828\times10^{-5} \times2\times10^{-4}}{500\times10^{-9}\times2}=8"